OpenStudy (5sos):

Pre-Cal Help! Will Upvote! and Fan!<3

OpenStudy (5sos):

http://prntscr.com/dd9zba

OpenStudy (tkhunny):

Encode the year. For 1991, 1991 - 1991 = 0, so use t = 0 for 1991 For 1995, 1995 - 1991 = 4, so use t = 4 for 1995 For 2011, 2011 - 1991 = 20, so use t = 20 for 2011 Your task is to find C(t) = 0.60 and then UNencode the result to a year. If you get C(q) = 0.60, then the proper year is q + 1991. It's all about the encoding.

OpenStudy (5sos):

Yeah this is all gibberish to me. XD sorri ;-; Hold on. I read the question and what does t mean?

OpenStudy (5sos):

Wait so I would plug in .60 for t and solve?

OpenStudy (tkhunny):

What to do is already there. Read very carefully. If you get C(q) = 0.60, then the proper year is q + 1991. Does that look like t = 0.60, like you asked, or something else? You need C = 0.60.

OpenStudy (5sos):

2007?

OpenStudy (tkhunny):

C(q) = 0.29(1.046)^{q} = 0.60 1.049^{q} = 0.60 / 0.29 = 2.069 q*log(1.049) = log(2.069) q = log(2.069) / log(1.049) = 0.31576 / 0.02078 = 15.195 1991 + 15.195 = 2006.2 It's possible, depending on how you are rounding, that 2007 is the correct answer. Please look at this response as an example of how to SHOW YOUR WORK.